Erica is 7 years older than Allen. In 5 years' time Erica will be twice as old as Allen. how old is Erica
Answer by ikleyn(52790)

.
E = A + 7           (1)     ("Erica is 7 years older than Allen.")
E + 5 = 2(A+5)      (2)     (" In 5 years . . . 


From eq(1), substitute the expression for E into eq(2). You will get


(A + 7) + 5 = 2*(A+5)


A + 12 = 2A + 10

12 - 10 = 2A - a

2 = A


Answer.  Allen is 2 years old;  Erica is 2 + 7 = 9 years old.

Solved.